If consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200. Y is domestic income and Y is private disposable income. Use the multiplayer application to export, to explain how a 100 billion decline in demand for export could affect the economie's GDP/income


1
Expert's answer
2021-05-16T17:48:05-0400

An export multiplier is a multiplier that applies to exports. The effect of a decline in demand by 100 billion on GDP/income will be computed by the export multiplier. Thus,

"M=\frac{change\space in\space import\space spending}{change\space in\space income}"

"=\frac{0.25}{0.5}"


"=0.5"

Export multiplier"=\frac{1}{(1-MPC)+M}"

"=\frac{1}{(1-0.75)+0.5}"


"=1.33"

Where (1 – MPC) is marginal propensity to save (MPS) and M is marginal propensity to import.

The export multiplier shows that the GDP/income will be decreased by 1.33 billion with the decline in the 100-billion decline in the demand.